OpenStudy (anonymous):

2c+6/4-c*4c-16/5c+15=

OpenStudy (debbieg):

Is this the expression? \(\Large 2c+\dfrac{6}{4}-c\cdot4c-\dfrac{16}{5c}+15=\) Because that's what you have written above. But I suspect what you MEAN is: \(\Large \dfrac{2c+6}{4-c}\cdot\dfrac{4c-16}{5c+15}=\) Which, if you don't want to use the equation editor, you can convey clearly with just a few well-placed parentheses: (2c+6)/(4-c)*(4c-16)/(5c+15)=

OpenStudy (anonymous):

yeah its the second one

OpenStudy (debbieg):

Assuming this: \(\Large \dfrac{2c+6}{4-c}\cdot\dfrac{4c-16}{5c+15}=\) Factor what you can (both num'rs, and the 2nd den'r, have GCF's that you can pull out). See what's left - something may cancel.

OpenStudy (debbieg):

Reduce fully first, then see what you have left.

OpenStudy (debbieg):

That is, factor each of: 2c+6 4c-16 5c+15 See what you have that might cancel.